OpenStudy (anonymous):

Question 1 Solve. 4/7 = x+3/21 A) x=9 B) x=15 C) x= 33 3/4 D) x=91 - Would it be A?? - I really need to pass this test! (procrastinated the whole break and now in a big hurry to raise my grade).

OpenStudy (bohotness):

n 4/7 = x+3/21 // - x+3/21 4/7-x-(3/21) = 0 4/7-x-1/7 = 0 3/7-x = 0 // - 3/7 -x = -3/7 // * -1

OpenStudy (bohotness):

4/7 = 21/x-3 multiply each side by x-3 (x-3) * 4/7 = 21 distribute (4x-12)/7 =21 4x - 12 = 147 4x = 159 x = 159/4 or 39.75 another way, flip each fraction over 7/4 = (x-3)/21 147/4 = x-3 147 / 4 = x - (12/4) 159/4 = x

OpenStudy (anonymous):

So the answer would be C? Sorry i am so tired. You are a big help.

OpenStudy (bohotness):

i mean its okay

OpenStudy (anonymous):

Im confused because 39.75 isn't one of the options ?

OpenStudy (bohotness):

4/7 = 12/21 12/21 = x + 3/21

OpenStudy (bohotness):

sorry about the other post this one might make it easy for you to pick

OpenStudy (anonymous):

Okay and what about this one. I just really need to get a good score. What is 24% of 75? A) 15 B) 16.5 C) 18 D) 22.5

OpenStudy (bohotness):

which one do you think is the answer?

OpenStudy (anonymous):

D) 22.5 for #2 right??? And for #1 I'm going with B is that correct?

OpenStudy (bohotness):

1 is a and 2 try again

OpenStudy (bohotness):

you close on 2

OpenStudy (anonymous):

18 for 2?

OpenStudy (bohotness):

yes

OpenStudy (bohotness):

do you have any more questions?

OpenStudy (anonymous):

and #3 is 36% of what number is 180? Is the answer 500??

OpenStudy (bohotness):

no try again lower its not in the 100s

OpenStudy (anonymous):

But there are no lower options?? :/ #4 : David drove 168 miles on 7 gallons of gasoline. What equation can be used to find how far can he can travel on a full tank of 16 gallons? Would it be 7/168 = 16/x ?

OpenStudy (bohotness):

3 Let the number be: x : 36%ofx=180 Or : 36100×x=180

OpenStudy (anonymous):

500 515 540 550 These are the only options for #3. Which would it be???

OpenStudy (bohotness):

what is the opion for 3?

OpenStudy (anonymous):

I went with 500.. but you said no? So is it 515, 540, or 550?

OpenStudy (anonymous):

I divided 180 by 0.36 and got 500. ?

OpenStudy (bohotness):

3 the anwser is 64.8

OpenStudy (bohotness):

i mean on 3 your answer is 64.8

OpenStudy (anonymous):

Thats what i thought! but theres no option for that i don't know what to do lol

OpenStudy (anonymous):

Ok whatever ill just skip that one and we can move on to number 4?

OpenStudy (anonymous):

thanks so much for the help by the way :) !

OpenStudy (bohotness):

and 4 is b

OpenStudy (bohotness):

7/168=16/x

OpenStudy (anonymous):

Ok thanks but i didn't give options for #4?